# From Single to Multi: How LLMs Hallucinate in Multi-Document Summarization

### v1

@@ Appendix C Model Access Details
We use the official OpenAI API [^7] https://platform.openai.com/docs/quickstart to access three LLMs, namely gpt-3.5-turbo-0125, gpt-4o-2024-05-13, and gpt-4o-mini-2024-07-18. To access gemini-1.5-flash we use Google’s Generative AI API [^8] https://ai.google.dev/gemini-api/docs/quickstart?authuser=2&lang=python. The remaining models, which include llama-v3p1-70b-instruct and qwen2-72b-instruct, are accessed through the Fireworks API [^9] https://fireworks.ai/models. Early on we also used the Fireworks API to evaluate LLMs, including Llama-3 (70B) (llama-v3-70b-instruct), and Mixtral (mixtral-8x7b-instruct-hf) but found numerous artifacts in their generations and, for that reason, excluded them from our analysis. For instance, Llama-3 (70B) generates preambles before generating list items (e.g., Here is a summary …, Here are the insights regarding …) but also postambles in 43% of its generations. In the postambles Llama-3 (70B) either re-iterate the instructions (e.g., Note: These insights are summarized from both Article 1 and Article 2, These insights … and are summarized in a concise and clear manner.) or summarizes the listed bullet points (e.g., These insights highlight the issues with Twitter’s complex tweet recommendation algorithm …, which is a critical step in improving the platform’s user experience and rebuilding trust with its users.). Conversely, mixtral-8x7b-instruct-hf’s generations either end abruptly, leading to incomplete summaries (e.g., Foot Locker aims to ensure that approximately 40% of its revenue, Twitter, under Elon Musk’s leadership, has announced plans to), or conclude with the total word count (e.g., (Word count: 241), (Note: The word count is exactly 200 words without the headings.)). Throughout our experiments, we explicitly focus on the evaluation of larger models due to their superior instruction following capability. All experiments were carried in 2024, with the main experiments (in Section 4) being conducted between July 1st to September 5th, and the mitigation experiments (described in Section 6) being conducted between September 10th and October 12th.
@@ Appendix D Automatic Metric Validation
In the main paper, we adopt an LLM-as-a-judge approach (Zheng et al., 2023) to determine whether the predicted insights faithfully cover the information in the input. Our evaluation approach is based on the metric proposed in the SummHay’s paper (Laban et al., 2024), which comprehensively compared the accuracy and human-alignment of different LLMs as evaluators, including gpt-4o (OpenAI, 2024b), Gemini-1.5-pro (Anil et al., 2024), and Claude (Anthropic, 2024). They find that among models they evaluate, Gemini-1.5-pro and gpt-4o were found to be positively correlated with human-level annotations. In this paper, however, we adopt gpt-4o-mini-2024-07-18 as a more cost-efficient but still competitive evaluator.101010gpt-4o-mini-2024-07-18 has been shown to be competitive with gpt-4o across various math, coding, and instruction-following benchmarks alike (OpenAI, 2024a; Chiang et al., 2024). To ensure that replacing the evaluator model does not negatively impact evaluation quality (Laban et al., 2024), we investigate the agreement between the two LLMs across 45.3k reference insights, spanning 4 summarization models and 2 distinct prompts. Given the complex nature of the evaluation process, our analysis is decomposed in terms of linking agreement—focused on whether models agree on which predicted insight (if any) covers the information of the reference insight—and coverage agreement—measuring to which extent the models agree on the coverage label (NO, PARTIAL, FULL).
@@
gpt-4o gpt-4o-mini Counts Frequency
@@ Appendix E Additional Results
  • Appendix E.6 studies the relationship between hallucinations and input documents;
@@ E.2 Analysis of the Number of Insights
Figure 18 shows the ratio of predicted-to-reference insights for the summaries generated for the two evaluated domains—news and conversation—as the number of input documents increases. By observing this ratio we can examine each models’ propensity to over- or under-generate insights. The first observation is that model behavior varies considerably depending on the prompt setting (“subtopic” or “subtopic+shared”) and the domain. Secondly, we observe that all models tend to over-estimate the number of relevant insights in the input documents when processing lower number of documents (N<3). In general, we find that
@@ E.3 F1-score Results
Figure 21 reports the F1-scores obtained for both “subtopic” and “subtopic+shared” settings across the news (Figure 21(a)) and conversation domain (Figure 21(b)).
@@ E.5 Correctness vs Outputs
E.6 Which documents do errors stem from?
So far, we have presented empirical evidence that models consistently hallucinate, irrespective of the number of input documents or task focus. In this section, we attempt to shed light on the source of model hallucinations by examining how they correlate with the input documents, i.e., we ask the question where do hallucinations come from? [^11] Prior work has studied this question with respect the faithfulness score (Huang et al., 2024). To answer this question, we first need to determine which input document (if any) the hallucinated insight is copied from.
Literature on data contamination.
To examine the correlation between errors and input documents it is necessary to determine which document (if any) the errors come from. A similar question has been researched in the context of LLMs and data contamination and memorization (Carlini et al., 2023; Xu et al., 2024b; Deng et al., 2024). To the best of our knowledge, there is no single best strategy to determine whether a test example is part of the pretraining data. Numerous techniques have been proposed, including those based on n-gram  (Brown et al., 2020; Chowdhery et al., 2024; Touvron et al., 2023) or sub-string matching (Carlini et al., 2023; Achiam et al., 2024), perplexity (Xu et al., 2024b), or by incorporating bloom filters into a fuzzy algorithm (Magnusson et al., 2023; Soldaini et al., 2024). For a more comprehensive discussion of the literature on data contamination, we refer the reader to the survey paper by Xu et al. (2024a).
Methodology.
We propose to modify the string matching contamination approach used to determine contamination in the GPT-4 paper (OpenAI, 2024b) and leave the exploration of other data contamination techniques for future work. Specifically, we assume that an hallucinated insight $i$ originates from a document $d$ if any 50-character string from $i$ matches a string in $d$.
Results.
We find that 2 out of 5 models show a near-uniform distribution of hallucinations across input documents (Figure 24), with gpt-3.5-turbo-0125 and gemini-1.5-flash exhibiting a modest recency bias towards the last documents, which are the closest to the list of instructions. Despite smaller in magnitude, this patterns also seems to be present in llama-v3p1-70b-instruct. Another factor influencing the results could be the models’ positional encoding, whose exploration we leave for future work.
Figure 24: Likelihood of an insight coming from a document (y-axis) based on its position in the input (x-axis) in the 10-combination setting. GPT-3.5-Turbo, Llama 3.1 (70B) and Gemini (Flash) exhibit recency bias, as their mistakes seem to be more likely to originate from later documents on average than from earlier ones.
@@ Appendix F Hallucination Taxonomy’s Annotation Details
To create the hallucination taxonomy, we collected human annotations for over 150 LLM-generated summaries and developed a taxonomy based on recurring mistakes. Due to the challenge of analyzing long documents, as noted in prior work (Chang et al., 2024), we limited our analysis to N=2, leaving the exploration of larger combinations for future research.
@@ G.1 Methods
  • Truncate summaries (Top-K): this method is based on our findings that LLMs’ insights generated earlier in the summary are more likely to be accurate than those generated later (see 4.4). Given a summary composed of $I$ insights (expressed as bullet-points), this method truncates the summary, keeping the first $K$ insights.121212 Early on, we also experimented with the removal of the first $L$ insights but found it to cause significant drops in recall and, therefore, chose not to include those results in the experiments.

### v2

@@ Appendix C Model Access Details
We use the official OpenAI API [^9] https://platform.openai.com/docs/quickstart to access three LLMs, namely gpt-3.5-turbo-0125, gpt-4o-2024-05-13, and gpt-4o-mini-2024-07-18. To access gemini-1.5-flash we use Google’s Generative AI API [^10] https://ai.google.dev/gemini-api/docs/quickstart?authuser=2&lang=python. The remaining models, which include llama-v3p1-70b-instruct and qwen2-72b-instruct, are accessed through the Fireworks API [^11] https://fireworks.ai/models. Early on we also used the Fireworks API to evaluate LLMs, including Llama-3 (70B) (llama-v3-70b-instruct), and Mixtral (mixtral-8x7b-instruct-hf) but found numerous artifacts in their generations and, for that reason, excluded them from our analysis. For instance, Llama-3 (70B) generates preambles before generating list items (e.g., Here is a summary …, Here are the insights regarding …) but also postambles in 43% of its generations. In the postambles Llama-3 (70B) either re-iterate the instructions (e.g., Note: These insights are summarized from both Article 1 and Article 2, These insights … and are summarized in a concise and clear manner.) or summarizes the listed bullet-points (e.g., These insights highlight the issues with Twitter’s complex tweet recommendation algorithm …, which is a critical step in improving the platform’s user experience and rebuilding trust with its users.). Conversely, mixtral-8x7b-instruct-hf’s generations either end abruptly, leading to incomplete summaries (e.g., Foot Locker aims to ensure that approximately 40% of its revenue, Twitter, under Elon Musk’s leadership, has announced plans to), or conclude with the total word count (e.g., (Word count: 241), (Note: The word count is exactly 200 words without the headings.)). Throughout our experiments, we explicitly focus on the evaluation of larger models due to their superior instruction following capability. All experiments were carried in 2024, with the main experiments (in Section 4) being conducted between July 1st to September 5th, and the mitigation experiments (described in Section 6) being conducted between September 10th and October 12th.
@@ Appendix D Automatic Metric Validation
In the main paper, we adopt an LLM-as-a-judge approach (Zheng et al., 2023) to determine whether the predicted insights faithfully cover the information in the input. Our evaluation approach is based on the metric proposed in the SummHay’s paper (Laban et al., 2024), which comprehensively compared the accuracy and human-alignment of different LLMs as evaluators, including gpt-4o (OpenAI, 2024b), Gemini-1.5-pro (Anil et al., 2024), and Claude (Anthropic, 2024). They find that among models they evaluate, Gemini-1.5-pro and gpt-4o were found to be positively correlated with human-level annotations. In this paper, however, we adopt gpt-4o-mini-2024-07-18 as a more cost-efficient but still competitive evaluator.121212gpt-4o-mini-2024-07-18 has been shown to be competitive with gpt-4o across various math, coding, and instruction-following benchmarks alike (OpenAI, 2024a; Chiang et al., 2024). To ensure that replacing the evaluator model does not negatively impact evaluation quality (Laban et al., 2024), we investigate the agreement between the two LLMs across 45.3k reference insights, spanning 4 summarization models and 2 distinct prompts. Given the complex nature of the evaluation process, our analysis is decomposed in terms of linking agreement—focused on whether models agree on which predicted insight (if any) covers the information of the reference insight—and coverage agreement—measuring to which extent the models agree on the coverage label (NO, PARTIAL, FULL).
@@
gpt-4o gpt-4o-mini Counts Relative Freq
@@ E.2 Analysis of the Number of Insights
Figure 19 shows the ratio of predicted-to-reference insights for the summaries generated for the two evaluated domains—news and conversation—as the number of input documents increases. By observing this ratio we can examine each models’ propensity to over- or under-generate insights. The first observation is that model behavior varies considerably depending on the prompt setting (“subtopic” or “subtopic+trustworthy”) and the domain. Secondly, we observe that all models tend to over-estimate the number of relevant insights in the input documents when processing lower number of documents (N<3). In general, we find that
@@ E.3 F1-score Results
Figure 22 reports the F1-scores obtained for both “subtopic” and “subtopic+trustworthy” settings across the news (Figure 22(a)) and conversation domain (Figure 22(b)).
@@ Appendix F Hallucination Taxonomy’s Annotation Details
To create the hallucination taxonomy, we collected human annotations for over 150 LLM-generated summaries and developed a taxonomy based on recurring mistakes. Due to the challenge of analyzing long documents, as noted in prior work (Chang et al., 2023), we limited our analysis to N=2, leaving the exploration of larger combinations for future research.
@@ G.1 Methods
  • Truncate summaries (Top-K): this method is based on our findings that LLMs’ insights generated earlier in the summary are more likely to be accurate than those generated later (see 4.5). Given a summary composed of $I$ insights (expressed as bullet-points), this method truncates the summary, keeping the first $K$ insights.131313 Early on, we also experimented with the removal of the first $L$ insights but found it to cause significant drops in recall and, therefore, chose not to include those results in the experiments.